OpenStudy (anonymous):

PLEASE HELP ASAP!! The arc corresponding to a central angle of 35 in a circle of radius 10 feet measures _____ feet. Round your answer to two decimal places.

OpenStudy (kainui):

 So if you know the distance around an entire circle compared to the angle it travels, then you have a proportion that you can compare like this:

OpenStudy (kainui):

\[\frac{ C }{ \theta }=\frac{ 2\pi r }{ \theta }=\frac{ 2\pi 10 }{ 2\pi }=\frac{ arc length }{ \theta }\] So just plug in to this, 35 and solve\[\frac{ 2\pi 10 }{ 360 }=\frac{ arc length }{ 35 }\]

OpenStudy (anonymous):

So it would be .17??

OpenStudy (anonymous):

@Kainui